1 Two point charges of equal and opposite charge 10-8 C sit a distance of 20 cm apart. What is the strength of the force they exert on each other? It it attractive or repulsive? 2 If one of the point charges in problem 1 is sitting on a table and has a mass of 10 g, how close must the second charge be to lift the charge up off the table? 3 What is the electric field strength required to lift the charge of the table in problem 2?
Answer #1

here,

the magnitude of charge on each point charges , q = 10^-8 C

the sepration between charges , r = 20 cm = 0.2 m

the strength of force , F = K * q^2 /r^2

F = 9 * 10^9 * (10^-8)^2 /0.2^2

F = 2.25 * 10^-5 N

the strength of force is 2.25 * 10^-5 N